(30 marks) Consider a perfectly competitive market in the short run. Assume that market demand is P=1000 - Q. Denoting firm level quantity by q, assume TC=30q-10q2+q3 so that MC=30-20q+3q2.


a) If there are 10 identical firm in the industry in the short run, what is the market equilibrium price and quantity?

b) Do firms make a profit or loss in the short run, and how much are these profits/losses?

c) What is the equilibrium price in the long run?

d) What will be equilibrium profit in the long run?

e) How many firms will there be in the long run?


1
Expert's answer
2021-04-05T07:25:42-0400

a) In the short run P=MC:


"1000-q=30-20q+3q^2,""3q^2-19q-970=0."

This quadratic equation has two roots: "q_1=21.4" and "q_2=-15.1". Since quantity can't be negative the correct answer "q_E=21.4." Then, we can find the equilibrium price:


"P_E=1000-21.4=\$978.6"

b) The profit can be found as follows:


"\pi=TR-TC=P_Eq_E-30q+10q^2-q^3,""\pi=978.6\cdot21.4-30\cdot21.4+10\cdot(21.4)^2-(21.4)^3=\$15079."

Therefore, the firms make a profit of $15079.

(c) In the long run ATC=MC:


"30-10q+q^2=30-20q+3q^2,""2q^2-10q=0,""q=5."

Then, we can find the equlibrium price in the long run:


"P_E=1000-5=\$995."

(d) Let's find the equilibrium profit in the long run:


"\pi=TR-TC=P_Eq_E-30q+10q^2-q^3,""\pi=995\cdot5-30\cdot5+10\cdot(5)^2-(5)^3=\$4950."

(e) Let's write the inverse demand function:


"Q=1000-P."

Then, we can find the market quantity:


"Q=1000-P=1000-\$995=5."

Finally, we can find the number of firms:


"nq=Q,""n=\dfrac{Q}{q}=\dfrac{5}{5}=1."